# Nearly-Linear Time Seeded Extractors with Short Seeds

### v1

@@ 5.2 A Recursive Construction
For any constant $\eta>0$ there exists a constant $C>0$ such that the following holds. For any positive integers $n$, $k\leq n$, and any $\varepsilon>0$ such that $k\geq C\log(n/\varepsilon)$ there exists a strong $(k,\varepsilon)$-seeded extractor
@@
  • if $k\geq 2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$;
  • if $k<2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step that corresponds to finding primitive elements of $O(\log\log n)$ fields $\mathds{F}_{q}$ with orders $q\leq\operatorname{poly}(n/\varepsilon)$ powers of $2$.
In a nutshell, our construction behind Theorem 5.7 works by considering two cases. If $\varepsilon>Cn^{3}\cdot 2^{-k/\log k}$, then we instantiate the recursive approach of Srinivasan and Zuckerman [SZ99] appropriately. Otherwise, we apply the recursive approach of Guruswami, Umans, and Vadhan [GUV09].
@@ 5.2.1 The (extremely) low-error case
In this section, we consider the lower error case of Theorem 5.7 where $\varepsilon\leq Cn^{3}\cdot 2^{-k/\log k}$. We instantiate the recursive approach from [GUV09, Section 4.3.3] appropriately, and analyze its time complexity. Crucially, because of our upper bound on $\varepsilon$, we will only need to run $O(\log\log n)$ levels of their recursive approach.
In order to obtain the statement of Theorem 5.7 for output length $m\geq(1-\eta)k$ with $\eta$ an arbitrarily small constant, it suffices to achieve output length $m=\Omega(k)$ and then apply Lemma 2.22 a constant number of times depending only on $\eta$. Therefore, we focus on achieving output length $m=\Omega(k)$.
@@
Furthermore, $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step that corresponds to finding primitive elements of $O(\log\log n)$ fields $\mathds{F}_{q}$ with orders $q\leq\operatorname{poly}(n/\varepsilon)$, each a power of $2$.
@@
For the base case $i(k)=0$, which holds when $k\leq 8d$, we choose $\mathsf{Ext}_{0}$ to be the $(k,\varepsilon_{0})$-seeded extractor defined as follows. On input an $(n,k)$-source $X$,
@@
    Apply the lossy RS strong condenser $\mathsf{RSCond}$ (Theorem 3.11) on $X$, instantiated with $\alpha=1/400$ and error $\varepsilon^{\prime}_{0}=\varepsilon_{0}/2$. When $C$ is large enough we have $k\geq(2+\alpha)\log(1/\varepsilon^{\prime}_{0})$, and require a seed $Y_{1}$ of length $d_{1}\leq C_{0}\log(n/\varepsilon^{\prime}_{0})$, for some constant $C_{0}>0$. The corresponding output $X^{\prime}$ satisfies $Y_{1}\circ X^{\prime}\approx_{\varepsilon^{\prime}_{0}}Y_{1}\circ Z$, for some $(n^{\prime},k^{\prime})$-source $Z$ with $k^{\prime}\geq(1-2\alpha)n^{\prime}=(1-1/200)n^{\prime}$.
@@
    Let $\mathsf{Ext}^{\prime}_{0}\colon\{0,1\}^{n^{\prime}}\times\{0,1\}^{d_{2}}\to\{0,1\}^{m^{\prime}}$ be the average-case strong $(k^{\prime},\varepsilon^{\prime}_{0})$-seeded extractor from Lemma 2.25 instantiated with $t=10$, which requires a seed $Y_{2}$ of length $d_{2}\leq k^{\prime}/10+C^{\prime}_{0}\log(n^{\prime}/\varepsilon^{\prime}_{0})$ for some constant $C^{\prime}_{0}>0$ and has output length $m^{\prime}\geq k^{\prime}/2$. The conditions for the invocation of Lemma 2.25 with $t=10$ are satisfied since $k^{\prime}\geq(1-1/200)n^{\prime}=(1-\frac{1}{20t})n^{\prime}$ and
@@
We set $Y=Y_{1}\circ Y_{2}$ and define $\mathsf{Ext}_{0}(X,Y)=\mathsf{Ext}^{\prime}_{0}(\mathsf{RSCond}(X,Y_{1}),Y_{2})$. From the discussion above, we have
$Y\circ\mathsf{Ext}_{0}(X,Y)=Y_{1}\circ Y_{2}\circ\mathsf{Ext}^{\prime}_{0}(\mathsf{RSCond}(X,Y_{1}),Y_{2})\approx_{\varepsilon^{\prime}_{0}}Y_{1}\circ Y_{2}\circ\mathsf{Ext}^{\prime}_{0}(Z,Y_{2})\approx_{\varepsilon^{\prime}_{0}}Y_{1}\circ Y_{2}\circ U_{m^{\prime}}.$
@@
provided that $d=C\log(n/\varepsilon)$ with $C$ a sufficiently large constant. The output length of $\mathsf{Ext}_{0}$ is $m^{\prime}\geq k^{\prime}/2\geq k/3$, since $k^{\prime}\geq(1-1/200)k$. Finally, both steps above take time $\widetilde{O}(n)$, and so $\mathsf{Ext}_{0}$ can be computed in time $\widetilde{O}(n)$ after a $\operatorname{polylog}(1/\varepsilon)$ preprocessing step.
@@
When $i(k)>0$, we assume the existence of the desired average-case strong extractors $\mathsf{Ext}_{i(k^{\prime})}$ for all $i(k^{\prime})<i(k)$ as the induction hypothesis. More precisely, we assume that for all $k^{\prime}$ such that $i(k^{\prime})<i(k)$ there exists a family of average-case strong $(k^{\prime},\varepsilon_{i(k^{\prime})})$-seeded extractors $\mathsf{Ext}_{i(k^{\prime})}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}\to\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{k^{\prime}/3}$ parameterized by $n$ computable in time $\widetilde{O}(n)$ after a one-time preprocessing step. We proceed as follows on input an $(n,k)$-source $X$:
@@
    Apply the lossy RS strong $(k,k^{\prime},\varepsilon_{1}=\varepsilon_{0}^{2})$-condenser $\mathsf{RSCond}$ (Theorem 3.11) on $X$ with $\alpha=1/20$ and a seed $Y_{\mathsf{RS}}$ of length $d_{\mathsf{RS}}\leq C_{\mathsf{RS}}\log(n/\varepsilon_{0})$. Since $k>8d\geq\frac{\log(1/\varepsilon_{1})}{\alpha(1+2\alpha)}$ if $C$ is a large enough constant, by the second part of Theorem 3.11 we know that with probability at least $1-\varepsilon_{0}$ over the choice of $Y_{\mathsf{RS}}=y$ it holds that the corresponding condenser output $X^{\prime}$ is $\varepsilon_{0}$-close to some $(n^{\prime},k^{\prime})$-source $Z$ with $k^{\prime}\geq(1-2\alpha)n^{\prime}=0.9n^{\prime}$. For the sake of exposition, from here onwards we work under such a good choice of the seed $Y_{\mathsf{RS}}$, and we will add the $\varepsilon_{0}$ slack term to the final error.
@@
    Split $X^{\prime}=X^{\prime}_{1}\circ X^{\prime}_{2}$ with $|X^{\prime}_{1}|=|X^{\prime}_{2}|=n^{\prime}/2\triangleq n^{\prime\prime}$. By Lemma 2.24 instantiated with $n^{\prime}$ and $\Delta=0.1n^{\prime}$ and the fact that $X^{\prime}$ is $\varepsilon_{0}$-close to an $(n^{\prime},k^{\prime})$-source, we get that $X^{\prime}_{1}\circ X^{\prime}_{2}$ is $(\varepsilon_{\mathsf{RS}}+2\varepsilon_{0}=3\varepsilon_{0})$-close to an $((n^{\prime\prime},n^{\prime\prime}),k^{\prime\prime}/n^{\prime\prime})$-block-source $W_{1}\circ W_{2}$ with
    $k^{\prime\prime}\geq k^{\prime}/2-\Delta-\log(1/\varepsilon_{0})\geq 0.4n^{\prime}-\log(1/\varepsilon_{0})\geq k/3,$ (3)
@@
    Apply the lossy RS strong $(k^{\prime\prime},k^{\prime\prime\prime},\varepsilon_{1}=\varepsilon_{0}^{2})$-condenser $\mathsf{RSCond}^{\prime}$ (Theorem 3.11) to $X^{\prime}_{2}$ with $\alpha=1/800$ and a seed $Y^{\prime}_{\mathsf{RS}}$ of length at most $d^{\prime}_{\mathsf{RS}}=C^{\prime}_{\mathsf{RS}}\log(n^{\prime\prime}/\varepsilon_{1})\leq C^{\prime}_{\mathsf{RS}}\log(n/\varepsilon_{0})$. From Item 2 and the data-processing inequality, we know that
    $Y^{\prime}_{\mathsf{RS}}\circ X^{\prime}_{1}\circ X^{\prime\prime}_{2}=Y^{\prime}_{\mathsf{RS}}\circ X^{\prime}_{1}\circ\mathsf{RSCond}(X^{\prime}_{2},Y^{\prime}_{\mathsf{RS}})\approx_{3\varepsilon_{0}}Y^{\prime}_{\mathsf{RS}}\circ W_{1}\circ\mathsf{RSCond}(W_{2},Y^{\prime}_{\mathsf{RS}}).$ (4)
    Since $(W_{2}|W_{1}=w_{1})$ is a $k^{\prime\prime}$-source for any $w_{1}$ in the support of $W_{1}$, we conclude from Theorem 3.11 and Equation 4 that
    $Y^{\prime}_{\mathsf{RS}}\circ W_{1}\circ\mathsf{RSCond}(W_{2},Y^{\prime}_{\mathsf{RS}})\approx_{\varepsilon_{1}}Y^{\prime}_{\mathsf{RS}}\circ W_{1}\circ\widetilde{W_{2}},$
    where $\widetilde{W_{2}}\sim\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n^{\prime\prime\prime}}$ and $\mathbf{H}_{\infty}(Y^{\prime}_{\mathsf{RS}}\circ\widetilde{W_{2}}|W_{1}=w_{1})\geq k^{\prime\prime\prime}+d^{\prime}_{\mathsf{RS}}$ for all $w_{1}$ in the support of $W_{1}$, with $n^{\prime\prime\prime}\geq k^{\prime\prime}\geq k^{\prime\prime\prime}\geq(1-1/400)n^{\prime\prime\prime}$. This is a valid invocation since $k^{\prime\prime}\geq k/3>8d/3>d\geq\frac{\log(1/\varepsilon_{1})}{\alpha(1+2\alpha)}$ by Equation 3. Therefore, by the second part of Theorem 3.11, with probability at least $1-\varepsilon_{0}$ over the choice of of $Y^{\prime}_{\mathsf{RS}}=y^{\prime}$ we get that
    $(W_{1}\circ\widetilde{W_{2}}|Y^{\prime}_{\mathsf{RS}}=y^{\prime})\approx_{\varepsilon_{0}}W_{1}\circ W^{\prime}_{2},$ (5)
    where $W^{\prime}_{2}\sim\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n^{\prime\prime\prime}}$ satisfies $\mathbf{H}_{\infty}(W^{\prime}_{2}|W_{1}=w_{1})\geq k^{\prime\prime\prime}\geq(1-1/400)n^{\prime\prime\prime}$. Fix such a good fixing of $Y^{\prime}_{\mathsf{RS}}$ from now onwards. As before, we will account for the probability $\varepsilon_{0}$ of fixing a bad seed in the final extractor error. Then, by combining Equations 4 and 5 we get that $X^{\prime}_{1}\circ X^{\prime\prime}_{2}$ is $(\varepsilon_{\mathsf{BS}}=4\varepsilon_{0})$-close to an $((n^{\prime\prime},n^{\prime\prime\prime}),k^{\prime\prime},k^{\prime\prime\prime})$-block source.
@@
    We will now apply block source extraction to $X^{\prime}_{1}\circ X^{\prime\prime}_{2}$, which we recall is $(\varepsilon_{\mathsf{BS}}=4\varepsilon_{0})$-close to an $((n^{\prime\prime},n^{\prime\prime\prime}),k^{\prime\prime},k^{\prime\prime\prime})$-block source. We instantiate Lemma 2.23 with $\mathsf{Ext}_{2}$ being the strong extractor from Lemma 2.25 with source input length $n^{\prime\prime\prime}$, min-entropy requirement $k^{\prime\prime\prime}$, error $\varepsilon_{\mathsf{BExt}}=\varepsilon_{0}$, output length $d$, and $t=16$. This requires a seed of length $d_{\mathsf{BExt}}\leq d/16+C^{\prime}_{0}\log(n/\varepsilon_{0})$. This instantiation of Lemma 2.25 is valid since $k^{\prime\prime\prime}\geq(1-1/400)n^{\prime\prime\prime}>(1-\frac{1}{20t})n^{\prime\prime\prime}$ and
@@
Items 1, 2, 3 and 4 above yield a strong seeded extractor $\mathsf{Ext}^{\prime}_{i(k)}\colon\{0,1\}^{n}\times\{0,1\}^{d^{\prime}}\to\{0,1\}^{m^{\prime}}$ with min-entropy requirement $k$, error $\varepsilon^{\prime}=\varepsilon_{i(k/3)}+\varepsilon_{\mathsf{BExt}}+\varepsilon_{\mathsf{BS}}+2\varepsilon_{0}=\varepsilon_{i(k/3)}+7\varepsilon_{0}$ (where the $2\varepsilon_{0}$ term comes from the two fixings of the seeds in the two condensing steps in Items 1 and 3), seed length
@@
To conclude the definition of $\mathsf{Ext}_{i(k)}$, we need to increase the output length of $\mathsf{Ext}^{\prime}_{i(k)}$ from $k/9$ to $k/3$. To that end, we use Lemma 2.22. Applying Lemma 2.22 once with $\mathsf{Ext}_{1}=\mathsf{Ext}^{\prime}_{i(k_{1})}$ with $k_{1}=k$ and $\mathsf{Ext}_{2}=\mathsf{Ext}^{\prime}_{i(k_{2})}$ with $k_{2}=k-k/9-1=8k/9-1$ and $g=1$ yields a strong $(k,3\varepsilon^{\prime})$-seeded extractor $\mathsf{Ext}^{\prime\prime}_{i(k)}$ with output length $(k_{1}+k_{2})/9\geq k(1-(8/9)^{2})-1$ and seed length $2(d/16+C^{\prime}\log(n/\varepsilon_{0}))=d/8+2C^{\prime}\log(n/\varepsilon_{0})$. Applying Lemma 2.22 again with $\mathsf{Ext}_{1}=\mathsf{Ext}^{\prime\prime}_{i(k_{1})}$ for $k_{1}=k$ and $\mathsf{Ext}_{2}=\mathsf{Ext}^{\prime\prime}_{i(k_{2})}$ for $k_{2}=(8/9)^{2}k$ and $g=1$ yields a strong $(k,9\varepsilon^{\prime})$-seeded extractor with output length $m\geq k(1-(8/9)^{4})-1\geq k/3$ and seed length $2(d/8+2C^{\prime}\log(n/\varepsilon_{0}))=d/4+4C^{\prime}\log(n/\varepsilon_{0})\leq d$, which we set as $\mathsf{Ext}_{i(k)}$. This second invocation of Lemma 2.22 is also valid, since $k_{2}=(8/9)^{2}k=k-(k(1-(8/9)^{2})-1)-1=k_{1}-m_{1}-g$. Note that the error $\varepsilon_{i(k)}=9\varepsilon^{\prime}=9\varepsilon_{i(k/3)}+63\varepsilon_{0}$, as desired.

### v2

@@ 4.2 A Recursive Construction
  1. 1.
    if $k\geq 2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$ and $\varepsilon\geq 2^{-Cn^{0.1}}$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$;
  2. 2.
    if $k\geq 2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$ and $\varepsilon<2^{-Cn^{0.1}}$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step, corresponding to generating $O(\log^{*}\!n)$ primes $q\leq\operatorname{poly}(n/\varepsilon)$;
  3. 3.
    if $k<2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step, corresponding to generating $O(\log\log n)$ primes $q\leq\operatorname{poly}(n/\varepsilon)$ and a primitive element for each field $\mathds{F}_{q}$.
In a nutshell, our construction behind theorem˜4.7 works by considering two cases. If $\varepsilon>Cn^{3}\cdot 2^{-k/\log k}$, then we instantiate the recursive approach of Srinivasan and Zuckerman [SZ99] appropriately. Otherwise, we apply the recursive approach of Guruswami, Umans, and Vadhan [GUV09].
@@ 4.2.1 The (extremely) low-error case
In this section, we consider the lower error case of theorem˜4.7 where $\varepsilon\leq Cn^{3}\cdot 2^{-k/\log k}$. We instantiate the recursive approach from [GUV09, Section 4.3.3] appropriately, and analyze its time complexity. Crucially, because of our upper bound on $\varepsilon$, we will only need to run $O(\log\log n)$ levels of their recursive approach.
In order to obtain the statement of theorem˜4.7 for output length $m\geq(1-\eta)k$ with $\eta$ an arbitrarily small constant, it suffices to achieve output length $m=\Omega(k)$ and then apply lemma˜2.21 a constant number of times depending only on $\eta$. Therefore, we focus on achieving output length $m=\Omega(k)$.
@@
Furthermore, $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step that corresponds to finding primitive elements of $O(\log\log n)$ fields $\mathds{F}_{q}$ with prime orders $q\leq\operatorname{poly}(n/\varepsilon)$.
@@
For the base case $i(k)=0$, which holds when $k\leq 8d$, we choose $\mathsf{Ext}_{0}$ to be the $(k,\varepsilon_{0})$-seeded extractor defined below. Before we define and analyze it formally, we informally discuss how the extractor works. Recall that we are aiming for seed length $d$, which in this base case satisfies $d\geq k/8$, output length $m\geq k/3$, and nearly-linear time complexity. Roughly speaking, on input an $(n,k)$-source $X$ we first apply the fast RS strong condenser to obtain an output $X^{\prime}$ that is close to a source with high min-entropy rate. Then, we apply the fast hash-based extractor from lemma˜2.24, which can be made to require only seed length $\approx k/t$ for a large constant $t$, to $X^{\prime}$ with a fresh seed. More formally,
@@
    Apply the lossy RS strong condenser $\mathsf{RSCond}$ (theorem˜3.17) on $X$, instantiated with $\alpha=1/400$ and error $\varepsilon^{\prime}_{0}=\varepsilon_{0}/2$. This requires a seed $Y_{1}$ of length $d_{1}\leq C_{0}\log(n/\varepsilon^{\prime}_{0})$, for some constant $C_{0}>0$, and is a valid invocation since $k\geq C\log(n/\varepsilon^{\prime}_{0})$ for a sufficiently large constant $C>0$. The corresponding output $X^{\prime}$ satisfies $(Y_{1},X^{\prime})\approx_{\varepsilon^{\prime}_{0}}(Y_{1},Z)$, for some $(n^{\prime},k^{\prime})$-source $Z$ with $k^{\prime}\geq(1-2\alpha)n^{\prime}=(1-1/200)n^{\prime}$.
@@
    Let $\mathsf{Ext}^{\prime}_{0}\colon\{0,1\}^{n^{\prime}}\times\{0,1\}^{d_{2}}\to\{0,1\}^{m^{\prime}}$ be the average-case strong $(k^{\prime},\varepsilon^{\prime}_{0})$-seeded extractor from lemma˜2.24 instantiated with $t=10$, which requires a seed $Y_{2}$ of length $d_{2}\leq k^{\prime}/10+C^{\prime}_{0}\log(n^{\prime}/\varepsilon^{\prime}_{0})$ for some constant $C^{\prime}_{0}>0$ and has output length $m^{\prime}\geq k^{\prime}/2$. The conditions for the invocation of lemma˜2.24 with $t=10$ are satisfied since $k^{\prime}\geq(1-1/200)n^{\prime}=(1-\frac{1}{20t})n^{\prime}$ and
@@
We set $Y=(Y_{1},Y_{2})$ and define $\mathsf{Ext}_{0}(X,Y)=\mathsf{Ext}^{\prime}_{0}(\mathsf{RSCond}(X,Y_{1}),Y_{2})$. We now argue that $\mathsf{Ext}_{0}$ is an extractor with the desired properties. From the discussion above, we have
$\mathopen{}\mathclose{{\left(Y,\mathsf{Ext}_{0}(X,Y)}}\right)=\mathopen{}\mathclose{{\left(Y_{1},Y_{2},\mathsf{Ext}^{\prime}_{0}(\mathsf{RSCond}(X,Y_{1}),Y_{2}}}\right)\approx_{\varepsilon^{\prime}_{0}}\mathopen{}\mathclose{{\left(Y_{1},Y_{2},\mathsf{Ext}^{\prime}_{0}(Z,Y_{2})}}\right)\approx_{\varepsilon^{\prime}_{0}}(Y_{1},Y_{2},U_{m^{\prime}}).$
@@
provided that $d=C\log(n/\varepsilon)$ with $C$ a sufficiently large constant. The output length of $\mathsf{Ext}_{0}$ is $m^{\prime}\geq k^{\prime}/2\geq k/3$, since $k^{\prime}\geq(1-1/200)k$. Finally, both steps above take time $\widetilde{O}(n)$, and so $\mathsf{Ext}_{0}$ can be computed in time $\widetilde{O}(n)$ after a one-time preprocessing step.
@@
When $i(k)>0$, we assume the existence of the desired average-case strong extractors $\mathsf{Ext}_{i(k^{\prime})}$ for all $i(k^{\prime})<i(k)$ as the induction hypothesis. More precisely, we assume that for all $k^{\prime}$ such that $i(k^{\prime})<i(k)$ there exists a family of average-case strong $(k^{\prime},\varepsilon_{i(k^{\prime})})$-seeded extractors $\mathsf{Ext}_{i(k^{\prime})}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}\times\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}\to\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{k^{\prime}/3}$ parameterized by $n$ computable in time $\widetilde{O}(n)$ after a one-time preprocessing step.
Intuitively, we will first use these extractors to construct an extractor $\mathsf{Ext}^{\prime}_{i(k)}$ with all the desired properties (min-entropy requirement $k$, small error $\varepsilon^{\prime}=\varepsilon_{i(k/3)}+7\varepsilon_{0}$, small seed length $d^{\prime}\approx d/16$) except that the output length is not large enough. Specifically, we will obtain output length $m=k/9$, but would like to get output length $m\geq k/3$ for the induction step. Roughly speaking, we proceed as follows on input an $(n,k)$-source $X$. First, we apply the fast RS strong condenser to $X$ to obtain a source $X^{\prime}$ with high min-entropy rate. Then, we split $X^{\prime}$ in half to create a block source $(X^{\prime}_{1},X^{\prime}_{2})$ with two blocks. This split decreases the entropies of $X^{\prime}_{1}$ and $X^{\prime}_{2}$, so we apply the fast RS strong condenser to $X^{\prime}_{2}$ to replace this second block by a block $X^{\prime\prime}_{2}$ with high min-entropy rate. Finally, we perform block source extraction on $(X^{\prime}_{1},X^{\prime\prime}_{2})$. Recall that all we are missing after this is a sufficiently large output length. The output length can be boosted via standard techniques, at the expense of slightly larger error and seed length.
More formally,
@@
    Apply the lossy RS strong $(k,k^{\prime},\varepsilon_{1}=\varepsilon_{0}^{2})$-condenser $\mathsf{RSCond}$ (theorem˜3.17) on $X$ with $\alpha=1/20$ and a seed $Y_{\mathsf{RS}}$ of length $d_{\mathsf{RS}}\leq C_{\mathsf{RS}}\log(n/\varepsilon_{0})$. This is valid since $k>8d\geq C_{\alpha}\log(n/\varepsilon)$ with $C_{\alpha}$ the constant from theorem˜3.17 with $\alpha=1/20$ if the constant $C$ in the theorem statement is large enough. By the second part of theorem˜3.17 we get that with probability at least $1-\varepsilon_{0}$ over the choice of $Y_{\mathsf{RS}}=y$ it holds that the corresponding condenser output $X^{\prime}$ is $\varepsilon_{0}$-close to some $(n^{\prime},k^{\prime})$-source $Z$ with $k^{\prime}\geq(1-2\alpha)n^{\prime}=0.9n^{\prime}$. For the sake of exposition, from here onwards we work under such a good choice of the seed $Y_{\mathsf{RS}}$, and we will add the $\varepsilon_{0}$ slack term to the final error.
@@
    Let $(X^{\prime}_{1},X^{\prime}_{2})$ correspond to the first two blocks of $\lfloor n^{\prime}/2\rfloor\triangleq n^{\prime\prime}$ bits of $X^{\prime}$. By lemma˜2.23 instantiated with $n^{\prime\prime}$ and $\Delta=0.1n^{\prime}$ and the fact that $X^{\prime}$ is $\varepsilon_{0}$-close to an $(n^{\prime},k^{\prime})$-source, we get that $(X^{\prime}_{1},X^{\prime}_{2})$ is $(\varepsilon_{0}+2\varepsilon_{0}=3\varepsilon_{0})$-close to an exact $((n^{\prime\prime},n^{\prime\prime}),k^{\prime\prime}/n^{\prime\prime})$-block-source $(W_{1},W_{2})$ with
    $k^{\prime\prime}\geq n^{\prime\prime}-\Delta-\log(1/\varepsilon_{0})\geq 0.4n^{\prime}-\log(1/\varepsilon_{0})-1\geq k/3,$ (11)
@@
    Apply the lossy RS strong $(k^{\prime\prime},k^{\prime\prime\prime},\varepsilon_{1}=\varepsilon_{0}^{2})$-condenser $\mathsf{RSCond}^{\prime}$ (theorem˜3.17) to $X^{\prime}_{2}$ with $\alpha=1/800$ and a seed $Y^{\prime}_{\mathsf{RS}}$ of length at most $d^{\prime}_{\mathsf{RS}}=C^{\prime}_{\mathsf{RS}}\log(n^{\prime\prime}/\varepsilon_{1})\leq C^{\prime}_{\mathsf{RS}}\log(n/\varepsilon_{0})$ to get $X_{2}^{\prime\prime}$. From item˜2 and the data-processing inequality, we know that
    $(Y^{\prime}_{\mathsf{RS}},X^{\prime}_{1},X^{\prime\prime}_{2})=\mathopen{}\mathclose{{\left(Y^{\prime}_{\mathsf{RS}},X^{\prime}_{1},\mathsf{RSCond}(X^{\prime}_{2},Y^{\prime}_{\mathsf{RS}})}}\right)\approx_{3\varepsilon_{0}}\mathopen{}\mathclose{{\left(Y^{\prime}_{\mathsf{RS}},W_{1},\mathsf{RSCond}(W_{2},Y^{\prime}_{\mathsf{RS}})}}\right).$ (12)
    Since $(W_{2}|W_{1}=w_{1})$ is a $k^{\prime\prime}$-source for any $w_{1}$ in the support of $W_{1}$, we conclude from theorem˜3.17 and eq.˜12 that
[…]
    where $\widetilde{W_{2}}\sim\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n^{\prime\prime\prime}}$ and $\mathbf{H}_{\infty}(Y^{\prime}_{\mathsf{RS}},\widetilde{W_{2}}|W_{1}=w_{1})\geq k^{\prime\prime\prime}+d^{\prime}_{\mathsf{RS}}$ for all $w_{1}$ in the support of $W_{1}$, with $n^{\prime\prime\prime}\geq k^{\prime\prime}\geq k^{\prime\prime\prime}\geq(1-1/400)n^{\prime\prime\prime}$. This is a valid invocation since $k^{\prime\prime}\geq k/3>8d/3>C\log(n/\varepsilon)$ for a large enough constant $C>0$ by eq.˜11. Therefore, by the second part of theorem˜3.17, with probability at least $1-\varepsilon_{0}$ over the choice of $Y^{\prime}_{\mathsf{RS}}=y^{\prime}$ we get that
    $\mathopen{}\mathclose{{\left(W_{1},\widetilde{W_{2}}}}\right)|\mathopen{}\mathclose{{\left\{{Y^{\prime}_{\mathsf{RS}}=y^{\prime}}}}\right\}\approx_{\varepsilon_{0}}(W_{1},W^{\prime}_{2}),$ (13)
    where $W^{\prime}_{2}\sim\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n^{\prime\prime\prime}}$ satisfies $\mathbf{H}_{\infty}(W^{\prime}_{2}|W_{1}=w_{1})\geq k^{\prime\prime\prime}\geq(1-1/400)n^{\prime\prime\prime}$. Fix such a good fixing of $Y^{\prime}_{\mathsf{RS}}$ from now onwards. As before, we will account for the probability $\varepsilon_{0}$ of fixing a bad seed in the final extractor error. Then, by combining eqs.˜12 and 13 we get that $(X^{\prime}_{1},X^{\prime\prime}_{2})$ is $(\varepsilon_{\mathsf{BS}}=4\varepsilon_{0})$-close to an $((n^{\prime\prime},n^{\prime\prime\prime}),k^{\prime\prime},k^{\prime\prime\prime})$-block source.
@@
    We will now apply block source extraction to $(X^{\prime}_{1},X^{\prime\prime}_{2})$, which we recall is $(\varepsilon_{\mathsf{BS}}=4\varepsilon_{0})$-close to an exact $((n^{\prime\prime},n^{\prime\prime\prime}),k^{\prime\prime},k^{\prime\prime\prime})$-block source. We instantiate lemma˜2.22 with $\mathsf{Ext}_{2}$ being the strong extractor from lemma˜2.24 with source input length $n^{\prime\prime\prime}$, min-entropy requirement $k^{\prime\prime\prime}$, error $\varepsilon_{\mathsf{BExt}}=\varepsilon_{0}$, output length $d$, and $t=16$. This requires a seed of length $d_{\mathsf{BExt}}\leq d/16+C^{\prime}_{0}\log(n/\varepsilon_{0})$. This instantiation of lemma˜2.24 is valid since $k^{\prime\prime\prime}\geq(1-1/400)n^{\prime\prime\prime}>(1-\frac{1}{20t})n^{\prime\prime\prime}$ and
@@
items˜1, 2, 3 and 4 above yield a strong seeded extractor $\mathsf{Ext}^{\prime}_{i(k)}\colon\{0,1\}^{n}\times\{0,1\}^{d^{\prime}}\to\{0,1\}^{m^{\prime}}$ with min-entropy requirement $k$, error $\varepsilon^{\prime}=\varepsilon_{i(k/3)}+\varepsilon_{\mathsf{BExt}}+\varepsilon_{\mathsf{BS}}+2\varepsilon_{0}=\varepsilon_{i(k/3)}+7\varepsilon_{0}$ (where the $2\varepsilon_{0}$ term comes from the two fixings of the seeds in the two condensing steps in items˜1 and 3), seed length
@@
Boosting the output length of $\mathsf{Ext}^{\prime}_{i(k)}$.
To conclude the definition of $\mathsf{Ext}_{i(k)}$, we need to increase the output length of $\mathsf{Ext}^{\prime}_{i(k)}$ from $k/9$ to $k/3$. To that end, we use lemma˜2.21. Applying lemma˜2.21 once with $\mathsf{Ext}_{1}=\mathsf{Ext}^{\prime}_{i(k_{1})}$ with $k_{1}=k$ and $\mathsf{Ext}_{2}=\mathsf{Ext}^{\prime}_{i(k_{2})}$ with $k_{2}=k-k/9-1=8k/9-1$ and $g=1$ yields a strong $(k,3\varepsilon^{\prime})$-seeded extractor $\mathsf{Ext}^{\prime\prime}_{i(k)}$ with output length $(k_{1}+k_{2})/9\geq k(1-(8/9)^{2})-1$ and seed length $2(d/16+C^{\prime}\log(n/\varepsilon_{0}))=d/8+2C^{\prime}\log(n/\varepsilon_{0})$. Applying lemma˜2.21 again with $\mathsf{Ext}_{1}=\mathsf{Ext}^{\prime\prime}_{i(k_{1})}$ for $k_{1}=k$ and $\mathsf{Ext}_{2}=\mathsf{Ext}^{\prime\prime}_{i(k_{2})}$ for $k_{2}=(8/9)^{2}k$ and $g=1$ yields a strong $(k,9\varepsilon^{\prime})$-seeded extractor with output length $m\geq k(1-(8/9)^{4})-1\geq k/3$ and seed length $2(d/8+2C^{\prime}\log(n/\varepsilon_{0}))=d/4+4C^{\prime}\log(n/\varepsilon_{0})\leq d$, which we set as $\mathsf{Ext}_{i(k)}$. This second invocation of lemma˜2.21 is also valid, since $k_{2}=(8/9)^{2}k=k-(k(1-(8/9)^{2})-1)-1=k_{1}-m_{1}-g$. Note that the error $\varepsilon_{i(k)}=9\varepsilon^{\prime}=9\varepsilon_{i(k/3)}+63\varepsilon_{0}$, as desired.